OpenStudy (anonymous):

angle HGF=16x+4, angle EGF=100 degrees, angle HGE=3x+11. Solve for x

jhonyy9 (jhonyy9):

do you can drawing it please ?

jhonyy9 (jhonyy9):

this HGEF wan being a rectangle ?

jhonyy9 (jhonyy9):

so a quadrilateral sure - yes ?

OpenStudy (anonymous):

HGF is an obtuse angle and segment GE is the bisector, sorta

OpenStudy (anonymous):

H. / / / / G. ---------------- E. \ \ \ \ F.

jhonyy9 (jhonyy9):

so you know that angle HGF = 16x+4 angle EGF = 100 degree angle HGE = 3x+11 and what you know again that GE is bisector so from what result that angle HGE congruent angle FGE - yes ?

jhonyy9 (jhonyy9):

and you can writing again that HGF = EGF + HGE - yes ?

OpenStudy (anonymous):

EGF is 110..100 was a typo

jhonyy9 (jhonyy9):

16x+4 = 110 + 3x+11 13x =117 x=117/13 x=9

jhonyy9 (jhonyy9):

do you understand it now why and how you get this result ?